*Just the answers are fine*

23) An internal control system is designed to

all the choices are correct

ensure reliable accounting

protect assets.

promote compliance to company policies

24) Your company buys a machine for $190,000. The machine has a useful life of 8 years and a salvage value of $10,000. The machine is estimated to produce 750,000 units of product over its useful life. The machine produces 105,000 units in its first year, and 109,000 units in its second year. In the second year, depreciation expense using the the units-of-production method, compared to the straight-line method, would be

$4,200 more

$2,640 less

$3,660 less

$3,660 more

25) ABC Company uses perpetual inventory. Details during January, 2020 are below:

     UNITS    COST PER UNIT    TOTAL

Balance, Jan 1      200 units      $5.00    $1,000

Purchase Jan 15 100 units    $6.00    600

Sale Jan 20 120 units

The company sells units for $10 each. A physical count of merchandise inventory on January 31 reveals that there are 180 units on hand

If the company uses FIFO, what is the gross profit for the month?

$1,423

$600

$1,376

$1,200
Answer #1

23) An internal control system is designed to - all the choices are correct

24) Depreciation under staright line (2nd year) = (190000 - 10000) / 8 = 22,500

Depreciation under units of production method (2nd year) = (190000 - 10000) / 750000 *109000 = 26,160

Thus depreciation under units of production method is $3660 more when compared with depreciation under straight line method.

Thus answer is $3660 more

25) Since the company uses FIFO basis, all the 120 units sold is out of the opening balance of inventory.

Thus cost of units sold = 120 units * 5 per unit = $600

Sales = 120 units * 10 per unit = $1200

Gross profit = Sales - cost of goods sold = 1200 - 600 = $600

Thus answer is $600

Kindly Up vote